# MC-GPT: Empowering Vision-and-Language Navigation with Memory Map and Reasoning Chains

### v1

@@ 3.4. Navigation Chain of Thoughts module
Inspired by the ICL (brown2020language, ; min2022rethinking, ) and CoT prompting (wei2022chain, ; zhang2022automatic, ), we expect that allowing LLM to reference human navigation examples in analogous scenarios can enrich the diversity of navigation strategies, thereby more effectively realizing “case-by-case analysis” in navigation reasoning. By querying a similar example with a chain of thoughts, we can offer a demonstration for the LLM.
@@
Figure 4. Two steps in the Navigation Chain of Thoughts: CoT mining and CoT querying. CoT mining aims to construct an example set comprising diverse reasoning chains, while CoT querying aims to select a similar example with an appropriate reasoning chain for the LLM’s navigation.
@@
Example set. The training set of the VLN task comprises “instruction-path” pairs, where annotators record viewpoint selections while executing the instructions. This training set inherently encompasses diverse examples with varying navigation strategies. As mentioned in Section 1, instructions leading to similar destinations often necessitate similar navigation strategies. To distinguish different navigation strategies of the training set, we utilize the destination’s room type as a criterion and create an example set based on this criterion. Specifically, for each “instruction-path” pair, we extract the ground-truth room type of the destination from the MatterPort3D simulator (anderson2018vision, ). Among the pairs of the training set, there exist a total of 27 room types. We select one example per room type to compose the example set.
@@
CoT Mining. The example set solely consists of instructions and corresponding viewpoint selections (“instruction-path”), which may not provide sufficient insight for the LLM to comprehend the reasoning behind these selections. Due to the unavailability of annotator explanations and the high cost associated with human annotation, we employ an LLM to emulate annotator behavior and mine the reasoning chain of each example. As shown in the left side of Fig. 4, we utilize the role prompting technique (john2023art, ) to instruct the LLM to play the role of the annotator: we present the LLM with perception and decision-making aspects, prompting it to explain the reason behind each viewpoint selection step by step. After the CoT mining, we extend the annotations of the example set from “instruction-path” to “instruction-path-explanation,” thereby establishing an example set with chains of thoughts. More details of our prompt can be found in Sec. 3.5.
@@
CoT Querying. Ideally, we aim to query an example with the same room type of destination as the test instance. However, determining the room type of the instance’s destination requires additional reasoning. To streamline this process, we use a parser to extract the room-type nouns from the instance’s instruction. We consider the first room-type noun as the destination’s room type based on idiomatic expressions (e.g., “go to the kitchen to get milk”). Subsequently, as shown in the right side of Fig. 4, we use the Sentence-BERT (reimers2019sentence, ) to encode the room-type nouns of the instance and those from the example set. The example with the closest cosine distance to the instance is selected as the demonstration example. In cases where the instruction lacks room-type nouns, we directly compute the cosine distance between the instance’s instruction and the room-type nouns from the example set. Then the selected demonstration example with the chain of thoughts is directly integrated into the prompt for the LLM.
3.5. Prompt manager
@@
To aid the LLM in understanding the agent’s perception and making the appropriate decisions, we introduce a prompt manager designed to organize all kinds of information. The prompt is depicted in Fig. 5.
Figure 5. Our prompt manager integrates two types of prompts, the \$Prefix prompt and the \$Input prompt. The former is designed to illustrate basic reasoning rules, while the latter provides perceptual information. The prompt manager is employed for both CoT mining and navigation inference. Thus, the symbol $\ast$ denotes content exclusively relevant during CoT mining, whereas $\ast\ast$∗ ∗ means content exclusive to navigation inference.
@@
3.6. Action prediction
The LLM receives the prompt as input and produces predictions. Among these output items, viewpoint selection is critical. The selected viewpoint corresponds to one of the viewpoints stored in the Memory Topological Map, allowing the shortest-path planner to compute a path to the target viewpoint.
4. Experimental setup
4.1. Datasets and Evaluation
We conduct experiments on two datasets: REVERIE and R2R. REVERIE dataset comprises 21,702 instructions, with an average instruction length of 18 words. The agent is required not only to navigate to the target location but also to point out the referred object by selecting the appropriate object bounding box. Considering the costs associated with the OpenAI API, we choose the validation unseen dataset which comprises 3,521 instructions to conduct experiments. As far as we know, we are the first LLM-based method to conduct experiments on the entire validation unseen dataset. The R2R dataset contains 21,567 instructions with an average length of 29 words. R2R is a well-established VLN benchmark, which features sequential instructions without object localization requirements. We follow the settings of NavGPT (zhou2023navgpt, ) and MapGPT (chen2024mapgpt, ) to evaluate the performance of our proposed MC-GPT on 72 scenes of the R2R dataset. We randomly sample one path for each scene, with three instructions per path, resulting in a total of 216 instructions.
We utilize the pre-defined evaluation metrics of the VLN task. (1) Trajectory Length (TL). (2) Success Rate (SR). (3) Oracle Success Rate (OSR): the agent’s success rate of having a goal viewpoint along the trajectory. (4) Success rate weighted by Path Length (SPL): the agent’s navigation success rate weighted by the trajectory length. (5) Navigation Error (NE). (6) Remote Grounding Success rate (RGS): the agent’s success rate of pointing out the referred object. (7) Remote Grounding Success rate weighted by navigation Path Length (RGSPL): the agent’s success rate of pointing out the referred object weighted by the trajectory length.
4.2. Implementation Details
We utilize the ChatGPT (gpt-3.5-turbo-1106 with 16K context window) as our LLM. Note that we have also experimented with Gemini (team2023gemini, ) and Llama2 (touvron2023llama, ), both of which did not perform as well as GPT-3.5. Therefore, we choose to present the results with GPT-3.5. To understand the visual content of the viewpoint’s images, we use the Faster-RCNN (girshick2015fast, ; anderson2018bottom, ) pre-trained on the Visual Genome (krishna2017visual, ) dataset to detect the surrounding objects and use the InstructBLIP-13B (li-etal-2023-lavis, ) as our VLM to caption the images of viewpoints.
5. Results
5.1. Main Results
@@
As shown in Tab. 1, our method yields significant improvements over methods in the Training category across all metrics, showing its superior few-shot capability. Compared to the Pre-training category, however, our method exhibits noticeable performance gaps. Pre-training methods typically adopt a two-step process that involves pre-training and fine-tuning, which extensively leverages VLN datasets to tailor strategies to specific datasets and domain knowledge. While this generates strong performance, it comes at the cost of high training overhead and limited generalization to novel environments. In contrast, our proposed MC-GPT adopts a versatile LLM, enabling easy adaptation to diverse downstream tasks without the need for additional training.
In Tab. 2, we evaluate MC-GPT on 72 various scenes of the R2R dataset. MC-GPT surpasses NavGPT (zhou2023navgpt, ) and MapGPT (chen2024mapgpt, ) across all metrics. The comparison demonstrates our advantages in memory construction and navigation strategy: on one hand, compared to NavGPT which uses an additional LLM to generate navigation memories, our memory map is more concise, has less noise, and does not require calling LLM. On the other hand, since the MapGPT is also equipped with a map, the advantage of our method suggests that the use of a navigation chain of thoughts is indeed helpful.
Table 1. Performance comparison on REVERIE.
@@
Seq2Seq-SF (anderson2018vision, ) 4.20 8.07 2.84 11.07 2.16 1.63
RCM (wang2019reinforced, ) 9.29 14.23 6.97 11.98 4.89 3.89
SMNA (self, ) 8.15 11.28 6.44 9.07 4.54 3.61
FAST-Short (ke2019tactical, ) 10.08 20.48 6.17 29.7 6.24 3.97
FAST-MATTN (qi2020reverie, ) 14.4 28.20 7.19 45.28 7.84 4.67
Pre-training RecBERT (hong2021vln, ) 30.67 35.02 24.90 16.78 18.77 15.27
Airbert (guhur2021airbert, ) 27.89 34.51 21.88 18.71 18.23 14.18
DUET (chen2022think, ) 46.98 51.07 33.73 22.11 32.15 23.03
@@
Table 2. Performance comparison on 72 various scenes of the R2R dataset. “Exp#” refers to the number of GPT experts and “Dist” the utilization of distance information.
@@
NavGPT (zhou2023navgpt, ) GPT-3.5 2 w/ 16.7 26.4 8.02
MapGPT (chen2024mapgpt, ) GPT-3.5 1 w/o 19.4 29.6 8.48
@@
Figure 6. Visualization of a successful navigation task. Initially, the agent is presented with multiple navigable viewpoints, including a stairway leading to the upper floor and several doors. Utilizing the proposed Memory Topological Map and Navigation Chain of Thoughts, the agent steadily advances toward the lounge area and finally identifies the target object – the trash can.
5.2. Ablation Study
@@
Is the navigation chain of thoughts essential? Our navigation chain of thoughts is included in a demonstration example, so we illustrate the necessity of the navigation thought chain by comparing the ablation effects of the demonstration example. As shown in Tab. 3, when no example is provided, that is, the LLM’s prompt only contains basic rules (prefix prompt), the results indicate that the LLM possesses commendable abilities in rule comprehension and navigation reasoning. Then we consider whether providing random demonstration examples is helpful: we compare two methods for querying a demonstration example, one based on instruction similarity, and the other based on destination room type similarity. Surprisingly, when using the former measure, the navigation performance is similar to that without using examples. This may be due to the presence of many function words in the instructions, leading to significant differences between the matched examples and the test scenes, and making the examples less useful as references. However, when using destination room type as a similarity measure, the performance is significantly improved, indicating that examples with similar destination room types are more meaningful references. Note that a demonstration example should contain complete navigation steps, otherwise they are not very helpful for the agent. In short, an appropriate demonstration example with a complete chain of thoughts can guide the agent for better navigation.
Does it help if the demonstration example is close to the test instance? As shown in Tab. 4, we conduct a comparative experiment: under the same conditions of using destination room type similarity for example querying, the closer the similarity, the more similar the scene between the demonstration example and the test instance. As shown in Tab. 4, when the destination room type of the example is the ground truth (RT-GT), the most suitable demonstration example can be matched, which is more helpful for navigation. When using Python’s regular expressions to extract room type from instructions (RT-RE), there is a slight decrease in performance, but it is within an acceptable range. In order to be closer to reality and reduce prediction costs, we adopt the latter method.
Table 3. Comparison of various settings of Navigation Chain of Thoughts on REVERIE. “Demo” refers to the utilization of a demonstration example and “Query-sim” refers to the similarity measure of the querying. “Step” refers to the number of steps of a demonstration example.
@@
$\times$× $\times$× $\times$× 18.00 30.00 8.83 20.13
[…]
Table 4. Effect of accuracy of destination room type prediction on navigation performance. “RT-GT” refers to the ground-truth room type, and “RT-RE” refers to the room type predicted by using Python’s regular expressions.
@@
5.3. Case Study
@@
6. Conclusion

### v2

@@ III-D Navigation Chain of Thoughts module
Inspired by the ICL [16, 35] and CoT prompting [36, 37], we expect that allowing LLM to reference human navigation examples in analogous scenarios can enrich the diversity of navigation strategies, thereby more effectively realizing “case-by-case analysis” in navigation reasoning. By querying a similar example with a chain of thoughts, we can offer a demonstration for the LLM.
@@
Figure 4: Two steps in the Navigation Chain of Thoughts: CoT mining and CoT querying. CoT mining aims to construct an example set comprising diverse reasoning chains, while CoT querying aims to select a similar example with an appropriate reasoning chain for the LLM’s navigation.
@@
The training set of the VLN task comprises “instruction-path” pairs, where annotators record viewpoint selections while executing the instructions. This training set inherently encompasses diverse examples with varying navigation strategies. As mentioned in Section I, instructions leading to similar destinations often necessitate similar navigation strategies. To distinguish different navigation strategies of the training set, we utilize the destination’s room type as a criterion and create a subset of the training set $S$. Specifically, for each “instruction-path” pair, we extract the ground-truth room type of the destination from the MatterPort3D simulator [1]. Among the pairs of the training set, there exist a total of 27 room types. We select one example per room type to compose the sampled training set $S$.
@@
CoT Mining. Since $S$ solely consists of instructions and corresponding viewpoint selections (“instruction-path”), this may not provide sufficient insight for the LLM to comprehend the reasoning behind these selections. Due to the unavailability of annotator explanations and the high cost associated with human annotation, we employ an LLM to emulate annotator behavior and mine the reasoning chain of each example. The pseudo-code of CoT mining is presented in Algorithm 2. Most of the processes are similar to the Algorithm 1, except the prompt since we expect the LLM to explain the viewpoint selection rather than predict the viewpoint. For the prompt $P_{t}$, as shown in the left side of Fig. 4, we utilize the role prompting technique [46] to instruct the LLM to play the role of the annotator: we present the LLM with perception and decision-making aspects, prompting it to explain the reason behind each viewpoint selection step by step. More details of our prompt $P_{t}$ can be found in Sec. III-E. After the CoT mining, we extend sample set $S$ from “instruction-path” to “instruction-path-explanation,” thereby establishing an example set $D$ with chains of thoughts.
@@
CoT Querying. Ideally, we aim to query an example with the same room type of destination as the test instance. However, determining the room type of the instance’s destination requires additional reasoning. To streamline this process, we use a parser to extract the room-type nouns from the instance’s instruction. We consider the first room-type noun as the destination’s room type based on idiomatic expressions (e.g., “go to the kitchen to get milk”). Subsequently, as shown in the right side of Fig. 4, we use the Sentence-BERT [43] to encode the room-type nouns of the instance and those from the example set. The example with the closest cosine distance to the instance is selected as the demonstration example. In cases where the instruction lacks room-type nouns, we directly compute the cosine distance between the instance’s instruction and the room-type nouns from the example set. Then the selected demonstration example with the chain of thoughts is directly integrated into the prompt for the LLM.
Here we represent the CoT querying mathematically. Let the example set be denoted as $D={d_{1},d_{2},\ldots,d_{Q}}$, where $Q$ is the total number of examples. We calculate the cosine similarity between the instruction $I$ and each example $d_{i}$, then we select the example $d_{q}$ with the highest cosine similarity to $I$:
$d_{q}=\rm{F_{C}}(\emph{I},\emph{D})=\arg\max_{\emph{d}_{i}\in\emph{D}}\left(\frac{\emph{I}\cdot\emph{d}_{i}}{\|\emph{I}\|\|\emph{d}_{i}\|}\right)$ (5)
[…]
III-E Prompt manager
@@
To aid the LLM in understanding the agent’s perception and making the appropriate decisions, we introduce a prompt manager $F_{P}$ designed to organize all kinds of information. The prompt is depicted in Fig. 5.
[…]
@@
III-F Action prediction
The LLM receives the prompt as input and produces predictions. Among these output items, viewpoint selection is critical. The selected viewpoint corresponds to one of the viewpoints stored in the Memory Topological Map, allowing the shortest-path planner $F_{S}$ to compute a path to the target viewpoint.
IV Experimental setup
IV-A Datasets
REVERIE dataset comprises 21,702 instructions, with an average instruction length of 18 words. The agent is required not only to navigate to the target location but also to point out the referred object by selecting the appropriate object bounding box.
R2R dataset contains 21,567 instructions with an average length of 29 words. R2R is a well-established VLN benchmark, which features sequential instructions without object localization requirements.
For the R2R dataset, we follow the experimental settings of previous works such as NavGPT [23] and MapGPT [25], conducting experiments solely on a subset of Val Unseen split of the R2R dataset. This setting serves two purposes: 1) considering the API costs, running the entire VLN dataset is impractical; 2) since neither NavGPT, MapGPT, nor our method involves training, any data split (validation seen, validation unseen, or test) remains unseen relative to the navigator. For the REVERIE dataset, we conduct evaluations on the entire Val Unseen split. To the best of our knowledge, our approach is the first LLM-core method to conduct evaluations on the entire val unseen dataset for the REVERIE dataset.
IV-B Evaluation Metrics
We utilize the pre-defined evaluation metrics of the VLN task. (1) Trajectory Length (TL): the agent’s path length in meters. (2) Success Rate (SR): the agent’s success rate of stopping at a viewpoint where the referred object can be observed. (3) Oracle Success Rate (OSR): the agent’s success rate of having a viewpoint along the trajectory where the referred object can be observed. (4) Success rate weighted by Path Length (SPL): the agent’s navigation success rate weighted by the length of the trajectory length. (5) Navigation Error (NE): the agent’s mean navigation error in meters. (6) Remote Grounding Success rate (RGS): the agent’s success rate of pointing out the referred object. (7) Remote Grounding Success rate weighted by navigation Path Length (RGSPL): the agent’s success rate of pointing out the referred object weighted by the trajectory length.
IV-C Implementation Details
We utilize the ChatGPT (gpt-3.5-turbo-1106 with 16K context window) and GPT-4o as our LLMs. Note that we have also experimented with Gemini [51] and Llama2 [18], both of which did not perform as well as GPT-3.5 and GPT-4o. Therefore, we choose to present the results with GPT-3.5 and GPT-4o. To understand the visual content of the viewpoint’s images, we use the Faster-RCNN [44, 38] pre-trained on the Visual Genome [45] dataset to detect the surrounding objects and use the InstructBLIP-13B [29] as our VLM to caption the images of viewpoints.
V Results
V-A Main Results
@@
As shown in Tab. I, our method yields significant improvements over methods in the Training category (which only fine-tune on the VLN datasets) across all metrics, demonstrating its superior capability in few-shot scenarios. However, compared to methods in the Pre-training category, our approach does show certain performance gaps. Pre-training methods typically involve a two-step process of pre-training followed by fine-tuning on VLN datasets. This process incurs high training overhead and often faces challenges in convergence. Additionally, these methods can suffer from limited generalization to novel environments. In contrast, our proposed MC-GPT leverages a versatile LLM, allowing for rapid adaptation to VLN tasks without the need for extensive additional training. Moreover, it enhances the model’s ability to generalize across different datasets and environments, making it a robust and efficient solution for varied applications in VLN tasks.
In Tab. II, we evaluate MC-GPT on 72 various scenes of the R2R dataset. By using GPT-3.5, our MC-GPT surpasses NavGPT [23] and MapGPT [25] across all metrics. The comparison demonstrates our advantages in memory construction and navigation strategy: on one hand, compared to NavGPT which uses an additional LLM to generate navigation memories, our memory map is more concise, has less noise, and does not require calling LLM. On the other hand, since the MapGPT is also equipped with a map, the advantage of our method suggests that the use of a navigation chain of thoughts is indeed helpful. We also conduct experiments with GPT-4o, and the comparison is shown in Tab. II. The results indicate that employing GPT-4o significantly enhances the navigator’s performance compared to using the GPT-3.5 model. Compared to the contemporary work MapGPT, our method, when using GPT-4o, shows slightly lower performance in Success Rate (SR) but surpasses MapGPT in both Oracle Success Rate (OSR) and Navigation Error (NE). This indicates that using the most advanced LLM can improve performance and complement the design presented in our paper. In other words, our method can demonstrate its effectiveness across different LLMs.
TABLE I: Performance comparison on REVERIE.
@@
Seq2Seq-SF [1] 4.20 8.07 2.84 11.07 2.16 1.63
RCM [6] 9.29 14.23 6.97 11.98 4.89 3.89
SMNA [48] 8.15 11.28 6.44 9.07 4.54 3.61
FAST-Short [49] 10.08 20.48 6.17 29.7 6.24 3.97
FAST-MATTN [2] 14.4 28.20 7.19 45.28 7.84 4.67
@@
TABLE II: Performance comparison on 72 various scenes of the R2R dataset. “Exp#” refers to the number of GPT experts and “Dist” the utilization of distance information.
@@
NavGPT [23] GPT-3.5 2 w/ 16.7 26.4 8.02
MapGPT [25] GPT-3.5 1 w/o 19.4 29.6 8.48
@@
MapGPT [25] GPT-4 1 w/o 41.2 61.6 5.80
MC-GPT (Ours) GPT-4o 1 w/o 32.1 68.8 5.42
Figure 6: Visualization of a successful navigation task. Initially, the agent is presented with multiple navigable viewpoints, including a stairway leading to the upper floor and several doors. Utilizing the proposed Memory Topological Map and Navigation Chain of Thoughts, the agent steadily advances toward the lounge area and finally identifies the target object – the trash can.
V-B Ablation Study
@@
Is the navigation chain of thoughts essential? Our navigation chain of thoughts is included in a demonstration example, so we illustrate the necessity of the navigation thought chain by comparing the ablation effects of the demonstration example. As shown in Tab. III, when no example is provided, that is, the LLM’s prompt only contains basic rules (prefix prompt), the results indicate that the LLM possesses commendable abilities in rule comprehension and navigation reasoning. Then we consider whether providing random demonstration examples is helpful: we compare two methods for querying a demonstration example, one based on instruction similarity, and the other based on destination room type similarity. Surprisingly, when using the former measure, the navigation performance is similar to that without using examples. This may be due to the presence of many function words in the instructions, leading to significant differences between the matched examples and the test scenes, and making the examples less useful as references. However, when using destination room type as a similarity measure, the performance is significantly improved, indicating that examples with similar destination room types are more meaningful references. Note that a demonstration example should contain complete navigation steps, otherwise they are not very helpful for the agent. In short, an appropriate demonstration example with a complete chain of thoughts can guide the agent for better navigation.
Does it help if the demonstration example is close to the test instance? As shown in Tab. IV, we conduct a comparative experiment: under the same conditions of using destination room type similarity for example querying, the closer the similarity, the more similar the scene between the demonstration example and the test instance. As shown in Tab. IV, when the destination room type of the example is the ground truth (RT-GT), the most suitable demonstration example can be matched, which is more helpful for navigation. When using Python’s regular expressions to extract room type from instructions (RT-RE), there is a slight decrease in performance, but it is within an acceptable range. In order to be closer to reality and reduce prediction costs, we adopt the latter method.
TABLE III: Comparison of various settings of Navigation Chain of Thoughts on REVERIE. “Demo” refers to the utilization of a demonstration example and “Query-sim” refers to the similarity measure of the querying. “Step” refers to the number of steps of a demonstration example.
@@
$\times$ $\times$ $\times$ 18.00 30.00 8.83 20.13
$\surd$ Instruction All 17.00 29.00 9.78 21.78
$\surd$ Room type 1 19.00 34.00 8.73 22.12
$\surd$ Room type All 23.00 34.00 11.16 23.80
TABLE IV: Effect of accuracy of destination room type prediction on navigation performance. “RT-GT” refers to the ground-truth room type, and “RT-RE” refers to the room type predicted by using Python’s regular expressions.
@@
Figure 7: Examples showcasing the conversion from visual observations to textual descriptions. InstructBLIP exhibits enhanced description capabilities leveraging our object detection context (a). Furthermore, in specific scenarios, it can articulate object orientations (b), identify room categories (c), and employ quantifiers (d) for descriptive purposes.
V-C Case Study
@@
V-D Cases of image captioning
In Section III-C, we use InstructBLIP-13B [29] for generating captions from observations. In this section, we present the specific prompt used with InstructBLIP-13B and provide several illustrative examples. The size of the input image is 640×480 pixels, containing the results of object detection. When objects are detected in the image, we prompt the InstructBLIP with the object detection results: “Context: This is a local view of a room with several objects including [detected objects]. Provide a brief description of the image”. When no objects are detected in the image, we simply prompt the InstructBLIP: “Provide a brief description of the image”.
In Fig. 7, we present some examples of our transformation from visual observations to textual descriptions. We find that InstructBLIP can better describe observations using our provided object detection results and, in some cases, can also describe object orientations, determine room types, and use quantifiers for descriptions. However, it is inevitable for InstructBLIP to generate illusions, which is highly related to the working mode of generative models. Therefore, relying solely on InstructBLIP’s perception is insufficient for the navigator. This is also one of the motivations behind proposing the memory topological map in this paper: to reduce noise and improve accuracy.
VI Conclusion